OpenStudy (anonymous):

Complete the standard form of the equation. -3y=8-4x _____x+_______y=8

OpenStudy (texaschic101):

add 4x to both sides and what do you get ?

OpenStudy (anonymous):

I'm not sure what x is so I don't understand it

OpenStudy (texaschic101):

we are not solving for x or y... -3y = 8 - 4x --- add 4x to both sides 4x - 3y = 8 - 4x + 4x....when we simplify, you will notice that the 4x's on the right side cancel....leaving you with : 4x - 3y = 8....which is the same as 4x + (-3y) = 8. Understand ?

OpenStudy (anonymous):

Yes thankyou

OpenStudy (texaschic101):

YW :)